Here are the relevant dates and interest rates:
Point A: August 1998: iff = 5.5%
Point B: September 1998: Lowered interest rates to iff =5.25%
Point C: October 1998: Lowered interest rates to iff = 5.00%
Point D: November 1998: Lowered interest rates to iff = 4.75%

Note importantly, we are modeling the behavior of the federal funds rate during this period. The forecasted reserve demand at this time is given below. For simplicity, this reserve demand function is stable (constant) throughout this exercise:
Rd = 950 - 110 iff

What is the value of the Reserve Supply in August 1998? Hint: use the Rd equation and the value of iff to get the value for Rd. Then remember that Rd=Rs in equilibrium!
Answer #1

Rd = 950 - 110 iff

August 1998 iff = 5.5%

Rd = 950 - 110 (5.5)

Rd = 950 - 605 = 345

Rs = Rd in Equilibrium. So, Rs = 345.

So, Value of Reserve Supply in August = 345.